# A Comprehensive Survey on Self-Supervised Learning for Recommendation

### v2

@@ 4.2.4. Adversarial Learning in SeqRec
  • Adversarial Learning Paradigm in SeqRec encompasses differential and non-differential methods. Differential methods, such as SAO (Manotumruksa and Yilmaz, 2020), SRecGAN (Lu et al., 2021), ACVAE (Xie et al., 2021), and ECGAN-Rec (Ni et al., 2023), utilize continuous operations for gradient propagation and avoid discrete sampling. SSRGAN (Lv et al., 2021) employs Gumble-Softmax for differentiable sampling to overcome gradient blocking. In non-differential methods, reinforcement learning optimizes the generator’s parameters after generating items through non-differentiable sampling and feeding them to the discriminator. AOS4Rec (Zhao et al., 2021a) employs the WGAN (Arjovsky et al., 2017) concept for simultaneous optimization and the actor-critic algorithm (Konda and Tsitsiklis, 1999) for stable training, while MFGAN (Ren et al., 2020) uses policy gradients to improve generator predictions using the discriminator’s score as a reward.
@@ 4.3. Social Recommendation
SEPT (Yu et al., 2021a) 2021 KDD Data-based Score-based InfoNCE-based
MHCN (Yu et al., 2021b) 2021 WWW Data-based Natural Explicit
@@
DSL (Wang et al., 2023c) 2023 IJCAI Model-based Natural Explicit
@@
HGCL (Chen et al., 2023) 2023 WSDM Model-based Natural InfoNCE-based
@@
CKER (Pan and Chen, 2021) 2022 Mathematics Model-based Natural InfoNCE-based
KGIC (Zou et al., 2022b) 2022 CIKM Data-based Score-based InfoNCE-based
@@
MCCLK (Zou et al., 2022a) 2022 SIGIR Model-based Natural InfoNCE-based
@@
ML-KGCL (Chen and Xie, 2023) 2023 DASFAA Model-based Natural InfoNCE-based
@@
KACL (Wang et al., 2023e) 2023 WSDM Model-based Natural InfoNCE-based
GL Method Venue Year Generative Learning Paradigm Generation Target
KGRec (Yang et al., 2023a) 2023 KDD Maksed Autoencoding Triplets in the KG
DiffKG (Jiang et al., 2024) 2024 WSDM Denoising Diffusion Triplets in the KG
@@ 4.3.2. Contrastive Learning in SocRec
  • View Creation in SocRec with CL can be categorized as data-based or model-based. The model-based approach emphasizes designing neural modules to encode features from social and collaborative data for contrastive views. Examples include KCGN (Huang et al., 2021) using behavior and temporal-aware networks, and SMIN (Long et al., 2021) and SDCRec (Du et al., 2022b) employing heterogeneous GNNs while considering social relationships. DSL (Wang et al., 2023c) utilizes GNNs for encoding distinct social-aware and collaborative-aware user embeddings. Conversely, data-based methods augment original graphs to generate diverse context-specific data. SEPT (Yu et al., 2021a) creates three graph views (preference, friend, and sharing) from user-item and user-user graphs, while MHCN (Yu et al., 2021b) generates hypergraphs among users using triangle motifs and applies hypergraph convolution. DcRec (Wu et al., 2022a) diversifies views through edge addition and node/edge dropout, and ReACL (Ji et al., 2023) constructs additional graphs using relationship-aware augmentation. In HGCL (Chen et al., 2023), a personalized knowledge transfer meta network encodes enhanced user and item representations for one view, while a lightweight GCN encodes another view for users/items.
  • Pair Sampling in SocRec is designed based on the encoded nature of the created views. In natural sampling methods, a user or item may have multiple views. For instance, KCGN and MHCN pair node embeddings with their encoded sub-graph or hypergraph embeddings, considering the local-global relationship. Furthermore, several methods (Long et al., 2021; Wu et al., 2022a; Wang et al., 2023c; Chen et al., 2023) encode multiple views for each node using graph augmentation or neural model design, treating the views of the same node as positive pairs. For score-based sampling, SEPT, SDCRec and ReACL label positive pairs by considering the representation similarly, and DUAL pre-computed a link-score for each edge in the graph based on the degree of the node, which is then utilized to choose positive pairs.
  • Contrastive Objective in these methods commonly relies on InfoNCE-based (Yu et al., 2021a; Wu et al., 2022a; Du et al., 2022b; Tao et al., 2022a; Ji et al., 2023; Chen et al., 2023) or JS-based objectives (Huang et al., 2021; Long et al., 2021). Notably, ReACL (Ji et al., 2023) introduces an Aug-InfoNCE contrastive loss, leveraging an expanded set of positive samples to improve consistency among similar nodes and enhance the framework’s generalization capability. In explicit optimization methods, MHCN employs a pair-wise ranking loss to maximize the mutual information between representations within a pair. On the other hand, DSL utilizes hinge loss to optimize the pair score, which proves beneficial in reducing the negative influence caused by noisy edges.
@@ 4.3.3. Adversarial Learning in SocRec
  • Adversarial Learning Paradigm. Adit et al. (Krishnan et al., 2019) proposed a modular adversarial approach to prevent model collapse in recommender systems with user-level social links. Their method utilized generative adversarial learning and a discriminator to differentiate sampled user pairs from both the interest and social domains. The optimization of parameters involved policy gradient due to the non-differentiable sampling operation. Similarly, DASO (Fan et al., 2019) leveraged the sampled user-user or user-item pairs from the generator and utilized policy gradient to optimize the parameters. RSGAN (Yu et al., 2019) uses the Gumbel-softmax technique for differentiable sampling and generates items based on social links. The discriminator in RSGAN aims to assign higher scores to both ground truth and socially aware items compared to negative items, following the social BPR optimization. In ESRF (Yu et al., 2020), a concrete selector layer is utilized to avoid non-differentiable sampling for neighbor users for discrimination.
  • Discrimination Target. In social recommendation with adversarial learning, the discrimination target aims to combine social information and collaborative information in a regularized manner. Adit et al.(Krishnan et al., 2019), RSGAN(Yu et al., 2019), and ESRF (Yu et al., 2020) all adopt a similar approach of obtaining user-user/item pairs through sampling from the generator. They then utilize the discriminator to enhance the quality of the generated samples. Notably, RSGAN (Yu et al., 2019) employs the generated items for discrimination, thereby improving the generator’s ability to generate more accurate items based on the observed ground truth items.
@@ 4.4.2. Contrastive Learning in KnoRec
  • View Creation in KnoRec involves the exploration of leveraging KGs in conjunction with the existing user-item graph to construct diverse perspectives. CKER (Pan and Chen, 2021) employs light-weight graph convolution and relation-aware convolution to encode interaction-aware and knowledge-aware item representations as two distinct views. KGIC (Zou et al., 2022b) utilizes the user-item interaction graph and KG to construct local and non-local graphs, encoding different views for nodes based on these two graphs. KGCL (Yang et al., 2022) and ML-KGCL (Chen and Xie, 2023) both employ graph augmentation, such as edge dropout, twice to create contrastive views. MCCLK (Zou et al., 2022a) first builds an item-item graph as the semantic view using relation-aware GNN, and then encodes collaborative and semantic representations for each node for contrastive learning. KRec-C2 (Peng et al., 2023) incorporates a category-level aggregation representation layer to obtain category-level signal features, which are contrasted with the original category features derived from the embedding function. HiCON (Wu et al., 2023a) leverages meta-paths to construct a high-order graph that encodes representations of users and items, contrasting them with representations obtained from the original interaction graph. KACL (Wang et al., 2023e) encodes the interaction view using an augmented user-item graph and KG-aware item representations based on the KG, providing another view for contrastive learning.
@@
  • Pair Sampling. Generally, in these methods, multiple views are created for each user and item node in the graph using various techniques (Zou et al., 2022b; Yang et al., 2022; Wang et al., 2023e). These views naturally form positive pairs for contrastive learning. In KGIC (Zou et al., 2022b), the distance between nodes on the local/non-local graph is used to select positive node pairs, enabling inter-graph and intra-graph interactive contrastive learning.
@@
  • Contrastive Objective. In all the mentioned methods, the contrastive objective is to optimize the mutual information between different views using the InfoNCE-based lower bound.
4.4.3. Generative Learning in KnoRec
@@
Generative learning also emerges as a novel SSL paradigm to further enhance the performance of KnoRec (Table 5). The intrinsic idea behind generative learning is to discover valuable knowledge triplets through generative tasks while mitigating the negative impact (Liu et al., 2020; Xiao et al., 2015) caused by noisy or irrelevant knowledge.
@@
  • Generative Learning Paradigm in KnoRec encompasses both masked autoencoding and denoising diffusion approaches. In KGRec (Yang et al., 2023a), a criterion score is initially computed for each knowledge triplet to identify the most valuable rationale triplets. Subsequently, the masking-and-reconstructing paradigm is employed to reconstruct these important triplets using node embeddings in the KG. This enables the encoding of information brought by these crucial triplets into the node representation, thereby enhancing recommendation performance. DiffKG (Jiang et al., 2024) employs the concept of denoising diffusion by adding Gaussian noise to the triplets in the KG and subsequently removing it. This approach effectively eliminates the inherent noise present in the knowledge graph, providing a cleaner and augmented KG that can be utilized for inference and recommendation purposes.
@@
  • Generation Target in both KGRec and DiffKG encompasses all the triplets present in the knowledge graph. This is because the original interactions within the KG may contain significant amounts of noise and redundant information that are irrelevant for recommendation (Yang et al., 2022). The generative learning process effectively generates valuable interactions within the KG through self-supervised learning, thereby achieving denoising.
@@ 4.5. Cross-domain Recommendation
PCRec (Wang et al., 2021) 2021 CogMI Data-based Natural InfoNCE-based
@@
SASS (Zhang et al., 2022c) 2022 CIKM Model-based Natural InfoNCE-based
@@
DR-MTCDR (Guo et al., 2023a) 2023 TOIS Data-based Natural InfoNCE-based
GL Method Venue Year Generative Learning Paradigm Generation Target
@@
VDEA (Liu et al., 2022c) 2022 SIGIR Variational Autoencoding User Rating to Items
@@
RecGURU (Li et al., 2022d) 2022 WSDM Differentiable Domain User Feature
@@
DA-DAN (Guo et al., 2023b) 2023 TOIS Differentiable Domain User Feature
@@ 4.5.2. Contrastive Learning in CroRec
  • View Creation. For data-based methods, PCRec (Wang et al., 2021) generates two distinct views using random walks on the source domain graph and encodes representations for contrastive pre-training. C2DSR (Cao et al., 2022a) creates single-domain and cross-domain sequences, pooling their prototype embeddings to form two positive views for contrastive learning. CLCDR (Chen et al., 2022c) replaces the BPR loss with a contrastive loss, creating views and positive pairs based on interaction data. DR-MTCDR (Guo et al., 2023a) employs node/edge dropping for graph augmentation, while CCDR (Xie et al., 2022a) uses a sub-graph-based data augmentation technique for view creation in its intra-domain CL. For model-based view creation, several methods consider representations learned by neural modules as distinct views. SASS (Zhang et al., 2022c) and CDRIB (Cao et al., 2022b) leverage contrastive loss to maximize differences between views of the same node, while CCDR (Xie et al., 2022a) uses neural-encoded representations of the same nodes from different domain data for inter-domain CL. SITN (Sun et al., 2023) encodes two user-specific sequences from source and target domains as distinct views and employs a clustering technique for an additional view. DCCDR (Zhang et al., 2023b) projects original representations into domain-specific and domain-invariant representations, maximizing mutual information between domain-invariant representations of users from two domains. For feature-based methods, CATCL (Xiao et al., 2023) adopts a feature-based view creation approach, similar to SimGCL, by adding noise to the node features during inference. This noise addition process generates diverse views that can be utilized for learning purposes.
  • Pair Sampling relies on natural sampling in these works, where multiple views are created for each data object. Positive pairs are formed by two views of the same object, while views from other objects serve as negative samples. Encoded representations from multiple views for the same user/item node can be paired, originating from different domains (Zhang et al., 2022c; Cao et al., 2022b; Xie et al., 2022a) or generated through data augmentation techniques (Wang et al., 2021; Xiao et al., 2023). Additionally, these representations can be encoded by specifically designed neural modules (Wang et al., 2021; Zhang et al., 2023b).
  • Contrastive Objective within CL-based cross-domain recommendation methods typically employs the InfoNCE-based objective to optimize the mutual information between positive views (Wang et al., 2021; Zhang et al., 2022c; Chen et al., 2022c; Xie et al., 2022a; Sun et al., 2023; Xiao et al., 2023; Zhang et al., 2023b; Guo et al., 2023a). Additionally, some methods utilize the JS-based lower bound for optimization. For instance, C2DSR (Cao et al., 2022a) corrupts the prototype representation to generate negative views, while CDRIB (Cao et al., 2022b) derives the JS-based contrastive term based on the information bottleneck regularization theory.
4.5.3. Generative Learning in CroRec

### v3

@@ 4.2.4. Adversarial Learning in SeqRec
  • Adversarial Learning Paradigm in SeqRec encompasses differential and non-differential methods. Differential methods, such as SAO (Manotumruksa et al., 2020), SRecGAN (Lu et al., 2021), ACVAE (Xie et al., 2021), and ECGAN-Rec (Ni et al., 2023), utilize continuous operations for gradient propagation and avoid discrete sampling. SSRGAN (Lv et al., 2021) employs Gumble-Softmax for differentiable sampling to overcome gradient blocking. In non-differential methods, reinforcement learning optimizes the generator’s parameters after generating items through non-differentiable sampling and feeding them to the discriminator. AOS4Rec (Zhao et al., 2021b) employs the WGAN (Arjovsky et al., 2017) concept for simultaneous optimization and the actor-critic algorithm (Konda et al., 1999) for stable training, while MFGAN (Ren et al., 2020) uses policy gradients to improve generator predictions using the discriminator’s score as a reward.
@@
4.2.5. In-depth Analysis of SSL-based Sequential Recommendation
SSL techniques in sequential recommendation each exhibit unique strengths and limitations tied to the temporal and contextual nature of sequential data. Contrastive learning methods, such as CL4SRec (Xie et al., 2022b) and CoSeRec (Liu et al., 2021d), excel at capturing nuanced sequential patterns through view augmentations and pairwise alignment, enhancing robustness against noise. However, their performance relies on the design of augmentation strategies and pair sampling, which may inadvertently distort temporal dependencies if not carefully regularized. Reconstructive approaches, exemplified by BERT4Rec (Sun et al., 2019) and DiffuASR (Liu et al., 2023a), leverage reconstruction objectives to model long-term dependencies via transformers or diffusion models, effectively addressing data sparsity through synthetic data generation. Adversarial learning, as seen in SRecGAN (Lu et al., 2021) and ECGAN-Rec (Ni et al., 2023), introduces adversarial training to improve recommendation diversity and robustness, but may struggle with instability due to the discrete nature of item sequences. Overall, different methods balance trade-offs between temporal coherence, data efficiency, and computational complexity inherent to sequential recommendation.
@@ 4.3. Social Recommendation
SEPT (Yu et al., 2021b) 2021 KDD Data-based Score-based InfoNCE-based
MHCN (Yu et al., 2021a) 2021 WWW Data-based Natural Explicit
@@
DSL (Wang et al., 2023d) 2023 IJCAI Model-based Natural Explicit
@@
HGCL (Chen et al., 2023b) 2023 WSDM Model-based Natural InfoNCE-based
@@
CKER (Pan et al., 2021) 2022 Mathematics Model-based Natural InfoNCE-based
KGIC (Zou et al., 2022a) 2022 CIKM Data-based Score-based InfoNCE-based
@@
MCCLK (Zou et al., 2022b) 2022 SIGIR Model-based Natural InfoNCE-based
@@
ML-KGCL (Chen et al., 2023a) 2023 DASFAA Model-based Natural InfoNCE-based
@@
KACL (Wang et al., 2023b) 2023 WSDM Model-based Natural InfoNCE-based
HEK-CL (Yuan et al., 2025) 2025 TOIS Data & Feature-based Natural Explicit
RL Method Venue Year Reconstructive Learning Paradigm Reconstruction Target
KGRec (Yang et al., 2023c) 2023 KDD Maksed Autoencoding Triplets in the KG
DiffKG (Jiang et al., 2024a) 2024 WSDM Denoising Diffusion Triplets in the KG
@@ 4.3.2. Contrastive Learning in SocRec
  • View Creation in SocRec with CL can be categorized as data-based or model-based. The model-based approach emphasizes designing neural modules to encode features from social and collaborative data for contrastive views. Examples include KCGN (Huang et al., 2021) using behavior and temporal-aware networks, and SMIN (Long et al., 2021) and SDCRec (Du et al., 2022b) employing heterogeneous GNNs while considering social relationships. DSL (Wang et al., 2023d) utilizes GNNs for encoding distinct social-aware and collaborative-aware user embeddings. Conversely, data-based methods augment original graphs to generate diverse context-specific data. SEPT (Yu et al., 2021b) creates three graph views (preference, friend, and sharing) from user-item and user-user graphs, while MHCN (Yu et al., 2021a) generates hypergraphs among users using triangle motifs and applies hypergraph convolution. DcRec (Wu et al., 2022a) diversifies views through edge addition and node/edge dropout, and ReACL (Ji et al., 2023) constructs additional graphs using relationship-aware augmentation. In HGCL (Chen et al., 2023b), a personalized knowledge transfer meta network encodes enhanced user and item representations for one view, while a lightweight GCN encodes another view for users/items.
  • Pair Sampling in SocRec is designed based on the encoded nature of the created views. In natural sampling methods, a user or item may have multiple views. For instance, KCGN and MHCN pair node embeddings with their encoded sub-graph or hypergraph embeddings, considering the local-global relationship. Furthermore, several methods (Long et al., 2021; Wu et al., 2022a; Wang et al., 2023d; Chen et al., 2023b) encode multiple views for each node using graph augmentation or neural model design, treating the views of the same node as positive pairs. For score-based sampling, SEPT, SDCRec and ReACL label positive pairs by considering the representation similarly, and DUAL pre-computed a link-score for each edge in the graph based on the degree of the node, which is then utilized to choose positive pairs.
  • Contrastive Objective in these methods commonly relies on InfoNCE-based (Yu et al., 2021b; Wu et al., 2022a; Du et al., 2022b; Tao et al., 2022a; Ji et al., 2023; Chen et al., 2023b) or JS-based objectives (Huang et al., 2021; Long et al., 2021). Notably, ReACL (Ji et al., 2023) introduces an Aug-InfoNCE contrastive loss, leveraging an expanded set of positive samples to improve consistency among similar nodes and enhance the framework’s generalization capability. In explicit optimization methods, MHCN employs a pair-wise ranking loss to maximize the mutual information between representations within a pair. On the other hand, DSL utilizes hinge loss to optimize the pair score, which proves beneficial in reducing the negative influence caused by noisy edges.
@@ 4.3.3. Adversarial Learning in SocRec
  • Adversarial Learning Paradigm in SocRec systems has emerged as a promising approach to leverage both user preferences and social relationships through generative adversarial frameworks. Adit et al. (Krishnan et al., 2019) proposed a modular adversarial approach to prevent model collapse in recommender systems with user-level social links, utilizing generative adversarial learning and a discriminator to differentiate sampled user pairs from interest and social domains. The parameter optimization involved policy gradient due to non-differentiable sampling operations. DASO (Fan et al., 2019) similarly leveraged sampled user-user or user-item pairs from the generator with policy gradient optimization. RSGAN (Yu et al., 2019) uses Gumbel-softmax for differentiable sampling and generates items based on social links, where the discriminator assigns higher scores to ground truth and socially aware items following social BPR optimization. In ESRF (Yu et al., 2020), a concrete selector layer is further utilized to avoid non-differentiable sampling for neighbor users for adversarial discrimination.
  • Discrimination Target in SocRec adversarial frameworks focuses on effectively integrating social and collaborative information through regularized discrimination mechanisms. Adit et al. (Krishnan et al., 2019), RSGAN (Yu et al., 2019), and ESRF (Yu et al., 2020) adopt similar approaches by sampling user-user/item pairs from generators and utilizing discriminators to enhance sample quality. Notably, RSGAN (Yu et al., 2019) specifically employs generated items for discrimination, improving the generator’s capability to produce more accurate items based on observed ground truth data.
4.3.4. In-depth Analysis of SSL-based Social Recommendation
In social recommendation, two primary techniques-contrastive learning, adversarial learning-exhibit distinct advantages and limitations. Contrastive learning, while effective in leveraging social relationships to enhance representation learning (Huang et al., 2021; Long et al., 2021), often requires complex view creation and pair sampling strategies that can introduce computational overhead and potential noise from social connections. Adversarial learning, on the other hand, excels in refining the quality of user interactions by employing generative models (Yu et al., 2019) to discern meaningful patterns in social data.
@@ 4.4.2. Contrastive Learning in KnoRec
  • View Creation in KnoRec involves the exploration of leveraging KGs in conjunction with the existing user-item graph to construct diverse perspectives. CKER (Pan et al., 2021) employs light-weight graph convolution and relation-aware convolution to encode interaction-aware and knowledge-aware item representations as two distinct views. KGIC (Zou et al., 2022a) utilizes the user-item interaction graph and KG to construct local and non-local graphs, encoding different views for nodes based on these two graphs. KGCL (Yang et al., 2022) and ML-KGCL (Chen et al., 2023a) both employ graph augmentation, such as edge dropout, twice to create contrastive views. MCCLK (Zou et al., 2022b) first builds an item-item graph as the semantic view using relation-aware GNN, and then encodes collaborative and semantic representations for each node for contrastive learning. KRec-C2 (Peng et al., 2023) incorporates a category-level aggregation representation layer to obtain category-level signal features, which are contrasted with the original category features derived from the embedding function. HiCON (Wu et al., 2023a) leverages meta-paths to construct a high-order graph that encodes representations of users and items, contrasting them with representations obtained from the original interaction graph. KACL (Wang et al., 2023b) encodes the interaction view using an augmented user-item graph and KG-aware item representations based on the KG, providing another view for contrastive learning.
@@
  • Pair Sampling in KnoRec contrastive learning methods focuses on constructing meaningful positive and negative sample pairs from knowledge graph-enhanced recommendation data. Generally, these methods create multiple views for each user and item node using various techniques (Zou et al., 2022a; Yang et al., 2022; Wang et al., 2023b), which naturally form positive pairs for contrastive learning. KGIC (Zou et al., 2022a) uses node distances on local/non-local graphs to select positive pairs, enabling inter-graph and intra-graph interactive contrastive learning.
@@
  • Contrastive Objective in KnoRec systems aims to maximize the mutual information between different knowledge-enhanced views through InfoNCE-based optimization. All mentioned methods optimize the mutual information between different views using the InfoNCE-based lower bound as their contrastive objective.
4.4.3. Reconstructive Learning in KnoRec
@@
Reconstructive learning also emerges as a novel paradigm to enhance the performance of KnoRec (Table 5). The intrinsic idea behind generative learning is to discover valuable knowledge triplets through generative tasks while mitigating the negative impact (Liu et al., 2020; Xiao et al., 2015) caused by noisy or irrelevant knowledge.
@@
  • Reconstructive Learning Paradigm in KnoRec encompasses both masked autoencoding and denoising diffusion approaches. In KGRec (Yang et al., 2023c), a criterion score is initially computed for each knowledge triplet to identify the most valuable rationale triplets. Subsequently, the masking-and-reconstructing paradigm is employed to reconstruct these important triplets using node embeddings in the KG. This enables the encoding of information brought by these crucial triplets into the node representation, thereby enhancing recommendation performance. DiffKG (Jiang et al., 2024a) employs the concept of denoising diffusion by adding Gaussian noise to the triplets in the KG and subsequently removing it. This approach effectively eliminates the inherent noise present in the knowledge graph, providing a cleaner and augmented KG that can be utilized for inference and recommendation purposes.
@@
  • Reconstruction Target in both KGRec and DiffKG encompasses all the triplets present in the knowledge graph. This is because the original interactions within the KG may contain significant amounts of noise and redundant information that are irrelevant for recommendation (Yang et al., 2022). The generative learning process effectively generates valuable interactions within the KG through self-supervised learning, thereby achieving denoising.
4.4.4. In-depth Analysis of SSL-based Knowledge-aware Recommendation
Knowledge-aware recommendation leverages SSL to address the challenges of sparse interactions and noisy knowledge integration. Contrastive learning methods, such as KGCL (Yang et al., 2022) and KGIC (Zou et al., 2022a), exploit multi-view representations derived from knowledge and interaction graphs, enhancing item/user embeddings by maximizing mutual information between views. Reconstructive approaches, like KGRec (Yang et al., 2023c) and DiffKG (Jiang et al., 2024a), focus on reconstructing or denoising knowledge triplets, implicitly filtering noise and reinforcing valuable structural patterns. Their performance relies on the quality of selected reconstruction targets, and care must be taken to avoid underperformance when critical triplets are sparse or ambiguous.
@@ 4.5. Cross-domain Recommendation
PCRec (Wang et al., 2021a) 2021 CogMI Data-based Natural InfoNCE-based
@@
SASS (Zhang et al., 2022d) 2022 CIKM Model-based Natural InfoNCE-based
@@
DR-MTCDR (Guo et al., 2023) 2023 TOIS Data-based Natural InfoNCE-based
RL Method Venue Year Reconstructive Learning Paradigm Reconstruction Target
@@
VDEA (Liu et al., 2022a) 2022 SIGIR Variational Autoencoding User Rating to Items
@@
RecGURU (Li et al., 2022b) 2022 WSDM Differentiable Domain User Feature
@@
DA-DAN (Guo and other, 2023) 2023 TOIS Differentiable Domain User Feature
@@ 4.5.2. Contrastive Learning in CroRec
  • View Creation in CroRec contrastive learning methods encompasses diverse strategies to construct meaningful representations across different domains through data-based, model-based, and feature-based approaches. For data-based methods, PCRec (Wang et al., 2021a) uses random walks on source domain graphs, C2DSR (Cao et al., 2022a) creates single-domain and cross-domain sequences with pooled prototype embeddings, CLCDR (Chen et al., 2022c) replaces BPR loss with contrastive loss based on interaction data, DR-MTCDR (Guo et al., 2023) employs node/edge dropping, and CCDR (Xie et al., 2022a) uses sub-graph augmentation for intra-domain CL. For model-based approaches, SASS (Zhang et al., 2022d) and CDRIB (Cao et al., 2022b) maximize differences between views of identical nodes, CCDR (Xie et al., 2022a) uses neural-encoded representations from different domains for inter-domain CL, SITN (Sun et al., 2023) encodes user sequences from both domains with clustering techniques, and DCCDR (Zhang et al., 2023b) projects representations into domain-specific and domain-invariant forms. For feature-based methods, CATCL (Xiao et al., 2023) adopts a feature-based view creation approach, similar to SimGCL, by adding noise to the node features during inference, which generates diverse views that can be utilized for learning purposes.
  • Pair Sampling relies on natural sampling in these works, where multiple views are created for each data object. Positive pairs are formed by two views of the same object, while views from other objects serve as negative samples. Encoded representations from multiple views for the same user/item node can be paired, originating from different domains (Zhang et al., 2022d; Cao et al., 2022b; Xie et al., 2022a) or generated through data augmentation techniques (Wang et al., 2021a; Xiao et al., 2023). Additionally, these representations can be encoded by specifically designed neural modules (Wang et al., 2021a; Zhang et al., 2023b).
  • Contrastive Objective within CL-based cross-domain recommendation methods typically employs the InfoNCE-based objective to optimize the mutual information between positive views (Wang et al., 2021a; Zhang et al., 2022d; Chen et al., 2022c; Xie et al., 2022a; Sun et al., 2023; Xiao et al., 2023; Zhang et al., 2023b; Guo et al., 2023). Additionally, some methods utilize the JS-based lower bound for optimization. For instance, C2DSR (Cao et al., 2022a) corrupts the prototype representation to generate negative views, while CDRIB (Cao et al., 2022b) derives the JS-based contrastive term based on the information bottleneck regularization theory.
4.5.3. Reconstructive Learning in CroRec